OpenStudy (18jonea):

When 2426 voters were polled, 25% said they were voting yes on an initiative measure.Find the margin of error and an interval that is likely to contain the true population proportion. ±4.9%; between 20.1% and 29.9% ±20%; between 5% and 45% ±49.3%; between 0.0% and 74.3% ±2%; between 23% and 27%

OpenStudy (18jonea):

@agent0smith

OpenStudy (agent0smith):

About the same as last time, show your work.

OpenStudy (18jonea):

im not sure on this one?

OpenStudy (agent0smith):

How is it different to the last? Hint: it's not.

OpenStudy (18jonea):

i know i just don't get this one

OpenStudy (agent0smith):

Do the same steps as the last one. I don't know what there is to not get?

OpenStudy (agent0smith):

If you could do that last one, you can do this one.

OpenStudy (18jonea):

i just don't know what numbers to put in

OpenStudy (agent0smith):

How were you able to do all this correctly: phat = 81/368 = 0.2201086957 ≈ 0.22 q = 1 - 0.2201086957 = 0.7798913043 = 0.78 P(0.22 - 1.96*sqrt(0.22*0.78/368) < π < 0.22 + 1.96*sqrt(0.22*0.78/368) ) = 0.95 P(0.22 - 0.04 < π < 0.22 + 0.04) = 0.95 P(0.18 < π < 0.26) = 0.95 sample proportion = 0.18 margin of error = 0.04 But be lost on this question?

jimthompson5910 (jim_thompson5910):

This problem doesn't state what the confidence level is. Does your teacher want you to assume 95% confidence?

OpenStudy (18jonea):

yeah i think so

OpenStudy (agent0smith):

Is that because you got that last one from here??? https://answers.yahoo.com/question/index?qid=20120623162446AAqJJMX

jimthompson5910 (jim_thompson5910):

You'll use this formula to find the margin of error (ME) \[\Large ME = 1.96*\sqrt{\frac{p*(1-p)}{n}}\] where we have 95% confidence, p is the sample proportion, n is the sample size

OpenStudy (18jonea):

no i had done this question for a previous assignment and just found my work for it

OpenStudy (agent0smith):

And it just happened to be copied and pasted from yahoo answers...

OpenStudy (18jonea):

p=When 2426 voters were polled, n=25%

OpenStudy (18jonea):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

p = 0.25 because 25% voted 'yes'

jimthompson5910 (jim_thompson5910):

n = 2426 is the sample size

OpenStudy (18jonea):

o ok

OpenStudy (18jonea):

sqrt.1875/2426= Me?

OpenStudy (18jonea):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how are you getting 1875?

jimthompson5910 (jim_thompson5910):

did you mean 0.25*0.75 = 0.1875 ??

OpenStudy (18jonea):

yes sorry

jimthompson5910 (jim_thompson5910):

use a calculator and tell me what `sqrt(0.1875/2426)` is equal to

OpenStudy (18jonea):

0.00879134326

OpenStudy (18jonea):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

now multiply that by 1.96 to get what?

OpenStudy (18jonea):

0.01723103279

jimthompson5910 (jim_thompson5910):

multiply that by 100 to convert to a percentage

OpenStudy (18jonea):

1.72310327997

jimthompson5910 (jim_thompson5910):

so roughly 1.7% which rounds to 2%

OpenStudy (18jonea):

right so answer d

jimthompson5910 (jim_thompson5910):

the confidence interval is then based on the margin of error 25% - 2% = 23% 25% + 2% = 27% the margin of error is from 23% to 27% most often in studies, you'll see "margin of error plus or minus 2%" or something like that

jimthompson5910 (jim_thompson5910):

`right so answer d` yes